# Approaching the thermodynamic limit of a bounded one-component plasma

### v1

@@ Abstract
The classical one-component plasma (OCP) bounded by a spherical surface reflecting ions (BOCP) is studied using molecular dynamics (MD). Simulations performed for a series of sufficiently large BOCP’s make it possible to establish the size dependencies for the investigated quantities and extrapolate them to the thermodynamic limit. In particular, the total electrostatic energy per ion is estimated in the limit of infinite BOCP in a wide range of the Coulomb coupling parameter $\Gamma$ from 0.03 to 1000 with the relative error of the order 0.1%. Calculated energies are by about 0.5% lower as compared to the modern Monte Carlo (MC) simulation data obtained by different authors at $\Gamma<30$ and almost coincide with the MC results at $\Gamma>175$. We introduce two more converging characteristic energies, the excess interatomic electrostatic energy and the excess ion–background electrostatic energy, which enable us to calculate the ionic compressibility factor inaccessible in conventional MC and MD simulation of the OCP with periodic boundary conditions. The derived wide-range ionic equation of state can be recommended for testing OCP simulations with various effective interaction potentials. Based on this equation, we propose an improved cutoff radius for the interionic forces implemented in LAMMPS and perform MD simulation of the OCP to demonstrate that location of the metastable region of the fluid–solid phase transition depends sensitively on this radius.
@@ I Introduction
A system of equal pointlike charges immersed in a uniform neutralizing background of fixed charge density is known as the one-component plasma (OCP) [1, 2, 3, 4]. This model is widely explored due to its simplicity that allows one to develop and test models for weakly and, especially, for strongly coupled plasmas. Introduction of a rigid background (although, some OCP modifications allow for the compressible background) is a convenient way to avoid instability of the Coulomb systems. In addition to an obvious importance of the OCP model for the theory, it has no less important applications in astrophysics because it is appropriate for the interiors of white dwarfs [5, 6, 7, 8, 9, 10], neutron stars [11], and for liquid alkali metals [12]. OCP turns out to be effective for such localized nonneutral plasma objects as trapped ions [13, 14] and the Coulomb clusters (Coulomb balls) in dusty plasmas [15, 16, 17, 18, 19] as well as for quasi-homogeneous dust plasma [20, 21]. Besides the classical OCP, quantum OCP is widely studied [22, 23], many more of them, in application to astrophysics [24, 25, 26]. The screened Coulomb (or Yukawa) systems are being studied no less intensively [27, 28, 29, 30, 31, 32]. In some studies, properties of the classical OCP are deduced in a limiting case of the Yukawa system with the infinite Debye length.
@@
For the classical OCP, most studied are the internal and Helmholtz free energy and the equation of state. Investigations are performed using analytical methods [33, 34, 35, 36, 37, 38, 32, 39], Monte Carlo (MC) [40, 41, 42, 43, 44, 45, 46, 47] and molecular dynamics (MD) simulation [48]. Extrapolation of the MC data to infinite number of ions makes it possible to obtain the most accurate (with the relative error about 0.1% and better) estimation of the system electrostatic energy. Extensively studied is melting of the OCP [49, 50, 51, 52, 53], the liquidlike to gaslike dynamical crossover [21, 54, 55]; the OCP transport properties are also of interest [55].
@@
Conventional conditions used in OCP simulations are periodic boundary conditions (PBC) that are inevitable in the case of long-range Coulomb forces. Since for such a system, electrostatic energy is represented as a series with conditional convergence, a method based on summation in reciprocal space developed by Ewald [56] (see also [57]) is used along with the modifications increasing its efficiency (e.g., [58, 59, 60]). For the PBC (or bulk) OCP, a single converging energy characteristic is the system total electrostatic energy per ion, at least, no other converging quantity can be obtained directly from the Ewald summation. In contrast, finite-size systems with the finite number of ions $N$ and an explicit boundary offer a possibility to search for such characteristics because each of them becomes formally finite and can be calculated directly by summation of individual Coulomb interactions, which does not require the Ewald summation. As is known [61], every Coulomb system reaches a thermodynamic limit at $N\to\infty$. Therefore, if it is possible to establish a size dependence for the quantity of interest or to demonstrate that it is almost size-independent at sufficiently large $N$ then a reliable extrapolation to the thermodynamic limit can be derived.
@@
In our recent study [62], we investigated large Coulomb clusters with a spherical neutralizing background at $\Gamma\gtrsim 200$ and demonstrated that for $N>2500$, a quasi-homogeneous cluster core can be found that can be treated as a satisfactory approach to the bulk system with compatible characteristics. We showed that the ionic compressibility factor $Z_{i}$ could be expressed in terms of the energies of interionic interaction and their interaction with the cluster background. However, we succeeded in a very crude estimation of $Z_{i}$ only at $\Gamma=500$ because the ion evaporation breaks noticeably both the system stability and its local quasineutrality at lower $\Gamma$.
@@
In this work, we consider the model of bounded OCP (BOCP) that is free from these shortcomings, in which a rigid boundary reflecting the ions specularly is situated at the surface of the background sphere. In this model, it is possible to investigate the entire range of $\Gamma$ from weakly to strongly coupled systems and to determine the size dependence for the quantities of interest. In addition to the conventional total electrostatic energy, we introduce two more energy quantities, the excess interatomic and ion–background electrostatic energy, and show them to be convergent in the thermodynamic limit. We also deduce the ionic pressure, i.e., the pressure of the ionic subsystem, as a linear combination of these energies. This enabled us to construct the ionic equation of state verified by determination of the ionic virial, which must hold for ordinary OCP.
@@
Yet another problem associated with the reciprocal space summation is an explicit dependence of the resulting interionic effective potential on the simulation cell length, which proves to be the potential decay length. This means that the Ewald potential, although decaying much more rapidly than the Coulomb one, is still the slowly decaying one. This is a source of ambiguity in taking into account the interaction of ions with their periodic images, which was noted already in the early study [40]. The most sensitive to this ambiguity is MD simulation, so that alternative methods for summation in the $k$-space were proposed like [58], in which the contribution from neighboring ions at the distance less than $r_{c}$ are included directly while that of distant ions are calculated by the $k$-space summation. Onegin et al. [63] first discovered that the virial determined in the MD simulation of OCP depends strongly on $r_{c}$, and moreover, this virial disagrees strongly with the pressure that is derived from the system potential energy. Again, the ab initio BOCP simulation does not encounter such problem, and we can compare the ionic pressure from our simulation with that for OCP. Thus, we propose to “fix” the MD simulations for OCP by taking $r_{c}$ as a function of $\Gamma$ when our $Z_{i}$ is used as reference data. As an illustration, we demonstrate that a different choice of $r_{c}$ changes drastically the width of metastable region of fluid–solid phase transition thus indicating the change in interfacial surface tension.
@@
In Sec. II.1, we introduce the main BOCP energy characteristics and explore them in limiting cases; in Sec. II.2, we derive the ionic equation of state as well as its excess ionic Helmholtz free energy. The method of MD simulation along with its justification and typical parameters is presented in Sec. III.1. In Sec. III.2, the MD results for BOCP are analyzed and discussed. Secs. IV.1 and IV.2 are devoted to the MD simulation of the OCP with PBC: the details of simulation method are discussed in Sec. IV.1; in Sec. IV.2, the recommended dependence $r_{c}(\Gamma)$ is presented, and the effect of this dependence on the kinetics of fluid–solid transition is illustrated.
@@ II.1 Energy characteristics of the bounded one-component plasma
We treat a system of $N$ particles (ions) with the mass $m$ and charge $ze$, where $e$ is the elementary charge, on a uniform background of the total charge zeN-zeN bounded by a sphere of the radius $R$. To prevent ion evaporation at high temperature, i.e., departure of the faster ions to infinity, the system is bounded by a rigid spherical boundary, from which the ions are reflected specularly. Since we develop an asymptotic model of the OCP, its radius $R_{b}$ must coincide with that of the background; otherwise, the system cannot be homogeneous even at low $\Gamma$. However, finiteness of the MD integration time step requires that $R_{b}$ is somewhat less than $R$ (Sec. III.1).
In what follows, we will term such system the BOCP. In contrast to OCP, here, another key parameter $N$ is added to the Coulomb coupling parameter $\Gamma=\epsilon/T$, where $\epsilon={z^{2}}{e^{2}}/{r_{s}}$ is the characteristic energy, ${r_{s}}={(3/4\pi{\bar{n}})^{1/3}}=R{N^{-1/3}}$ is the ion sphere radius, $\bar{n}=3N/4\pi{R^{3}}$ is the average ion number density, $T$ is the temperature in energy units; BOCP is assumed to be in the equilibrium state, and $N$ is sufficiently large. Hereafter, we will use the Coulomb units such as the length $r_{s}$, the energy $\epsilon$, and the time ${\tau_{0}}={r_{s}}{(m/\epsilon)^{1/2}}={(mr_{s}^{2}/\Gamma T)^{1/2}}$, which coincides with the inverse oscillation frequency of an ion in the ion sphere (in these units, $T=1/\Gamma$, $R=N^{1/3}$, and $\bar{n}=3/4\pi$).
@@
${u_{p}}=\displaystyle{1\over N}\sum\limits_{k<l}\displaystyle{1\over r_{kl}}=\displaystyle{1\over 2N}\iint\limits_{V_{R}}{\displaystyle{{n({\bf{r}}_{1}^{\prime})n({\bf{r}}_{2}^{\prime})}\over{{r^{\prime}_{12}}}}\,d{\bf{r}}_{1}^{\prime}d{\bf{r}}_{2}^{\prime}}$ (2)
is the interionic interaction energy per ion, $r_{12}=|{\bf{r}}_{1}-{\bf{r}}_{2}|$, $n({\bf{r}}^{\prime})=\sum\nolimits_{k=1}^{N}{\delta({\bf{r}}^{\prime}-{{\bf{r}}_{k}})}$ is the ion number density, ${\bf{r}}_{k}$ is the radius-vector of the $k$th ion, and the integration is performed in the background volume $V_{R}$;
@@
$={1\over{2N}}\int\limits_{{V_{R}}}{{{r^{\prime}}^{2}}n({\bf{r}^{\prime}})}\,d{\bf{r}^{\prime}}-{5\over 2}{u_{c}}$
@@
is the background energy per ion. Equation (3) implies that origin of the coordinate system coincides with the background center. Although $u$ converges at $N\to\infty$, $u_{p}$ and $u_{b}$ diverge as $N^{2/3}$; however, one can define renormalized quantities that will be termed the excess interatomic and ion–background electrostatic energy
@@
We will demonstrate convergence of $u_{p\mathrm{ex}}$ and $u_{b\mathrm{ex}}$ in the cases of weakly and strongly coupled systems. If the system is gaslike, and the ternary and higher ion correlations can be neglected, then
$\left\langle{n({{\bf{r}}_{1}})n({{\bf{r}}_{2}})}\right\rangle=\left\langle{n(0)n(r_{12})}\right\rangle={\left({\displaystyle{3\over{4\pi}}}\right)^{2}}{f_{p}}(r_{12}),$ (6)
where angular brackets denote time averaging, and
@@
is the radial distribution function and $N_{r}$ is the average number of ions within a sphere of the radius $r$ around an ion. Similarly, $\left\langle{n({\bf{r}})}\right\rangle=\left\langle{n(r)}\right\rangle=(3/4\pi){f_{c}}(r)$, where
@@
$\left\langle{{u_{p\mathrm{ex}}}}\right\rangle$ $\simeq{1\over 2N}{\left({\displaystyle{3\over{4\pi}}}\right)^{2}}\iint\limits_{V_{R}}{\displaystyle{{f_{p}}(r_{1}){f_{p}}(r_{2})-1\over{r_{12}}}\,d{\bf{r}}_{1}d{\bf{r}}_{2}}$ (9)
@@
$\left\langle{{u_{b\mathrm{ex}}}}\right\rangle\simeq\displaystyle{3\over{2N}}\int\limits_{0}^{R}{{r^{4}}\left[{{f_{c}}(r)-1}\right]\,dr}.$ (10)
@@
Thus, in the thermodynamic limit, we obtain an obvious result: all excess electrostatic energies of ideal gas vanish:
@@
where $u_{\infty}=\mathop{\lim}\limits_{N\to\infty}\left\langle{u}\right\rangle$. Note that if we treat separately the system of ions in the field of background then the system potential energy $u_{i}={u_{p}}+{u_{b}}$ diverges as $N\to\infty$. However, the excess potential energy $u_{i}+u_{c}={u_{p\mathrm{ex}}}+{u_{b\mathrm{ex}}}$ does converge.
Consider the case $\Gamma\to\infty$, in which the ion sphere model (see, e.g., [2]) proved to be most effective for the Coulomb clusters [64, 62]. This model implies that the entire system is divided into spherical overlapping spheres of the radius $r_{s}$, in which single ions reside, and all interparticle and particle–background interaction can be reduced to the interaction of an ion with the compensating background of its sphere. Thus, the net force acting on an ion is simply ${\bf{f}}=-{\bf{r}}$, where the origin of the coordinate system is shifted to the center of an ion sphere and we omit subscripts due to identity of the ions. Based on the virial theorem [65] and an obvious equality for a harmonic oscillator $\left\langle{{r^{2}}}\right\rangle=\left\langle{{v^{2}}}\right\rangle=3/\Gamma$, we can write the ion compressibility factor as [62]
${Z_{i}}=1+\displaystyle{\Gamma\over{3N}}\sum\limits_{i=1}^{N}{\left\langle{{\bf{rf}}}\right\rangle}=1-\displaystyle{\Gamma\over 3}\left\langle{{r^{2}}}\right\rangle=0,$ (13)
@@
We will use the virial theorem one more time to express the ion pressure and corresponding compressibility factor in terms of the ion subsystem potential and kinetic energy. However, if we assume that the charge neutralizing background of BOCP is fixed, $r_{s}=\mathrm{const}$, then the energy $u_{p}$ (2) is a homogeneous function of order 1-1, $u_{b}+5u_{c}/2$, of order 2 (Eq. (3)), and $u_{c}$, of order 0. Bearing in mind that if the function $\varphi$ is a sum of homogeneous functions of order $k$, $\varphi(x)=\sum\nolimits_{k}{{\psi_{k}}(x)}$, then
${{d\varphi}\over{dx}}=\sum\limits_{k}{\displaystyle{{k{\psi_{k}}(x)}\over x}},$ (14)
and one can modify the virial theorem for the ions in the field of background as
@@
Obviously, (18) hold in the thermodynamic limit $N\to\infty$. Thus, we have demonstrated that the quantities (5) have thermodynamic limits at low and high $\Gamma$, which is indicative of the fact that these limits exist at arbitrary $\Gamma$ defining the ionic pressure for the bulk OCP.
Since an objective of this paper is investigation of the OCP properties, the size dependence $u(\Gamma,N)$ (1) is of interest. As it was demonstrated in [62], all quantities derived from the BOCP energy characteristics depend on $N$ very weakly, therefore, we adopt $\ln N$ as an argument. The simplest and most general form of the size dependence is then the expansion in powers of shifted $\ln N$, where we retain solely two terms:
@@
Here, $c_{1}$ and $c_{2}$ are the parameters to be defined from fitting the MD simulation results (Sec. III.2).

### v2

@@ Abstract
The classical one-component plasma (OCP) bounded by a spherical surface reflecting ions (BOCP) is studied using molecular dynamics (MD). Simulations performed for a series of sufficiently large BOCPs make it possible to establish the size dependences for the investigated quantities and extrapolate them to the thermodynamic limit. In particular, the total electrostatic energy per ion is estimated in the limit of infinite BOCP in a wide range of the Coulomb coupling parameter $\Gamma$ from 0.03 to 1000 with a relative error of about 0.1%. The calculated energies are lower by nearly 0.5% than the modern Monte Carlo (MC) simulation data obtained by different authors at $\Gamma<30$ and almost coincide with the MC results at $\Gamma>175$. We introduce two more converging characteristic energies, the excess interatomic energy and the excess ion–background energy, which enable us to calculate the ionic compressibility factor inaccessible for conventional MC and MD simulations of the OCP with periodic boundary conditions. The derived wide-range ionic equation of state can be recommended for testing OCP simulations with various effective interaction potentials. Based on this equation, we propose an improved cutoff radius for the interionic forces implemented in LAMMPS. We demonstrate that the kinetic and interfacial quantities determined for the classical OCP may be ambiguous. To check this, we perform a simulation of the OCP metastable region and find that its width depends sensitively on the cutoff radius decreasing with the increase of the latter.
@@ I Introduction
A system of equal pointlike charges immersed in a uniform neutralizing background of fixed charge density is known as the one-component plasma (OCP) [1, 2, 3, 4]. This model is widely explored due to its simplicity that allows one to develop and test models for weakly and, especially, strongly coupled plasmas. Introduction of a rigid background (although some OCP modifications allow for a compressible background) is a convenient way to avoid instability of Coulomb systems. In addition to obvious theoretical importance of OCP models, they have no less important applications in astrophysics because they are appropriate for the interiors of white dwarfs [5, 6, 7, 8, 9, 10] and neutron stars [11]. Such models are also appropriate for liquid alkali metals [12]. Since the confinement field configuration in ion traps is not much different from a parabolic potential, the OCP model is used in the physics of such localized non-neutral plasma objects as ion Coulomb crystals, which are extensively studied both theoretically and experimentally [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. In these studies, a crystallized core was observed and simulated, e.g., in spherical plasmas with more than $2\times 10^{5}$ ions, the formation of bcc crystals occupying the inner quarter of the plasma diameter was detected [18].
Another application of the OCP model is the Coulomb cluster (Coulomb ball) in dusty plasmas [23, 24, 25, 26, 27] as well as the quasi-homogeneous dusty plasma [28, 29]. Besides the classical OCP, quantum OCP is widely studied [30, 31], especially for astrophysics [32, 33, 34]. The screened Coulomb (or Yukawa) systems are being studied no less intensively [35, 36, 37, 38, 39, 40, 41]. In some studies, properties of the classical OCP are deduced in the limiting case of the Yukawa system with infinite Debye length. In the theoretical studies of spherical Coulomb crystals [42, 43] performed for extremely low temperatures and the system sizes up to $1.2\times 10^{5}$, the cluster interior with crystal ordering surrounded by a few shells on the outside was found.
@@
For the classical OCP, the most studied quantities are the internal energy, the Helmholtz free energy, and the equation of state (EOS). Investigations are performed using analytical methods [44, 45, 46, 47, 48, 49, 41, 50], Monte Carlo (MC) [51, 52, 53, 54, 55, 56, 57, 58], and molecular dynamics (MD) simulation [59]. Extrapolation of the MC data to infinite number of ions allows one to obtain the most accurate estimate of the system electrostatic energy (with the relative error about 0.1% and better). Extensively studied is melting of the OCP [60, 61, 62, 63, 64], the liquidlike to gaslike dynamical crossover [29, 65, 66]; the OCP transport properties are also of interest [66].
@@
Conventional conditions used in OCP simulations are periodic boundary conditions (PBC) [67], which are inevitable in the case of long-range Coulomb forces because they substantially minimize a strong effect of the system boundaries. PBC imply that the simulated system consists of a cubic supercell of the edge length $L$ containing the ions along with an infinite lattice of its periodic images. Each image cell is obtained by translating the original supercell by sLsL along one or more coordinate axes, where $s=\pm 1,\pm 2,\pm 3,\ldots$ Thus, an ion in the supercell with the radius-vector $\mathbf{r}$} has image ions at positions ${\mathbf{r}}+{\mathbf{s}}L$, where $\mathbf{s}$ is a vector of integers, and $\mathbf{s}\neq 0$. For such a system, the electrostatic energy is represented as a conditionally converging series. Since this series converges very slowly, a method based on summation in reciprocal space developed by Ewald [68] (see also [69]) is used along with the modifications increasing its efficiency (e.g., [70, 71, 72]). For the PBC (or bulk) OCP, the only convergent energy characteristic is the total electrostatic energy per ion; at least, no other converging quantity can be obtained directly from the Ewald summation. In contrast, finite-size systems with the finite number of ions NN and an explicit boundary offer a possibility to search for such characteristics because each of them becomes formally finite and can be calculated directly by summation of individual Coulomb interactions, which does not require the Ewald summation. As is known, under certain conditions, the Coulomb system reaches a thermodynamic limit at infinite size. Therefore, if it is possible to establish a size dependence for the quantity of interest at sufficiently large NN, then a reliable extrapolation to the thermodynamic limit can be derived.
@@
In our recent study [73], we investigated large Coulomb clusters with a finite spherical neutralizing background and demonstrated that for $N>2500$, a quasi-homogeneous cluster core can be found that can be treated as a satisfactory approach to the bulk system with compatible characteristics. Here, our results are very close to those obtained in [43]. In particular, we have also found that the core has a polycrystalline structure represented by bcc, fcc, and hcp crystallites. In contrast to [42, 43], where the Coulomb coupling parameter $\Gamma$ was set to $10^{5}\mbox{--}10^{6}$, we considered much lower $\Gamma\gtrsim 200$, which corresponded to the region of cluster melting crossover, and investigated the dependence of the spherical shell structure and the core size on $\Gamma$. In addition, we showed that the ionic compressibility factor could be expressed in terms of the energies of interionic interaction and their interaction with the cluster background. However, we only obtained a crude estimate of $Z_{i}$ at $\Gamma=500$ and noticed that the ion evaporation breaks noticeably both the system stability and its local quasineutrality at $\Gamma<200$.
@@
In this work, we consider a bounded OCP (BOCP), which is free from these shortcomings, with a rigid boundary that reflects the ions specularly and is situated at the surface of the background sphere. In this model, it is possible to investigate the entire range of $\Gamma$ from weakly to strongly coupled systems and to determine the size dependence for the quantities of interest. In addition to the conventional total electrostatic energy, we introduce two more energy quantities, the excess interatomic and ion–background electrostatic energy, and show them to be convergent in the thermodynamic limit. We also deduce the ionic pressure, i.e., the pressure of the ionic subsystem, as a linear combination of these energies. This enables us to construct the ionic EOS verified by determination of the ionic virial, which must hold for ordinary OCP.
@@
Yet another problem associated with the reciprocal space summation is an explicit dependence of the resulting interionic effective potential on the simulation cell length, which proves to be the potential decay length. This means that the Ewald potential, although decaying much more rapidly than the Coulomb one, is still a slowly decaying one. This is a source of ambiguity in taking into account the interaction of ions with their periodic images, which was noted already in the early study [51]. The most sensitive to this ambiguity is MD simulation, so that alternative methods for summation in $k$-space were proposed such as that in [70], in which the contribution from neighboring ions at distances less than a properly chosen Coulomb cutoff radius $r_{c}$ is included directly while that of distant ions is calculated by the $k$-space summation. Onegin et al. [74] were the first who discovered that the virial determined in the MD simulation of OCP depends on $r_{c}$, and moreover, this virial disagrees strongly with the pressure that is derived from the system potential energy. Again, the ab initio BOCP simulation does not encounter such problem, and we can compare the ionic pressure from our simulation with that for OCP. Thus, we propose to “fix” the MD simulations for OCP by taking $r_{c}$ as a function of $\Gamma$ when our $Z_{i}$ is used as reference data. As an illustration, we demonstrate that a different choice of $r_{c}$ changes significantly the width of metastable region of fluid–solid phase transition thus indicating the change in interfacial surface tension.
@@
In Sec. II.1, we introduce the main BOCP energy characteristics and explore them in limiting cases; in Sec. II.2, we derive the ionic EOS as well as its excess ionic Helmholtz free energy. The method of MD simulation along with its justification and typical parameters is presented in Sec. III.1. In Sec. III.2, the MD results for BOCP are analyzed and discussed. Secs. IV.1 and IV.2 are devoted to the MD simulation of the OCP with PBC: the details of simulation method are discussed in Sec. IV.1; in Sec. IV.2, the recommended dependence $r_{c}(\Gamma)$ is presented. The effect of this dependence on the kinetics of fluid–solid phase transition is examined in Sec. V.
@@ II.1 Energy characteristics of the bounded one-component plasma
We treat a system of NN particles (ions) with the mass $m$ and charge zeze, where $e$ is the elementary charge, on a uniform background of the total charge zeN-zeN bounded by a sphere of the radius $R$. To prevent ion evaporation at high temperature, i.e., departure of the faster ions to infinity, the system is bounded by a rigid spherical boundary, from which the ions are reflected specularly. Since we develop an asymptotic model of the OCP, its radius $R_{b}$ must coincide with that of the background; otherwise, the system cannot be homogeneous even at low $\Gamma$. However, finiteness of the MD integration time step requires that $R_{b}$ is somewhat less than $R$ (Sec. III.1).
In what follows, we will term such system the BOCP. In contrast to OCP, here, another key parameter NN is added to the Coulomb coupling parameter $\Gamma=\epsilon/T$, where $\epsilon={z^{2}}{e^{2}}/{r_{s}}$ is the characteristic electrostatic energy per ion, ${r_{s}}={(3/4\pi{\bar{n}})^{1/3}}=R{N^{-1/3}}$ is the ion sphere radius, $\bar{n}=3N/4\pi{R^{3}}$ is the average ion number density, $T$ is the temperature in energy units; BOCP is assumed to be in the equilibrium state, and NN is sufficiently large. Hereafter, we use the Coulomb units by setting $r_{s}=m=\epsilon=ze=1$. Then the unit time is ${\tau_{0}}={r_{s}}{(m/\epsilon)^{1/2}}={(mr_{s}^{2}/\Gamma T)^{1/2}}$, which coincides with the inverse oscillation frequency of an ion in the ion sphere. In the Coulomb units, $T=1/\Gamma$, $R=N^{1/3}$, and $\bar{n}=3/4\pi$.
@@
${u_{p}}=\displaystyle{1\over N}\sum\limits_{k<l}\displaystyle{1\over r_{kl}}=\displaystyle{1\over 2N}\iint\limits_{V_{R}}{\displaystyle{{n({\mathbf{r}}_{1}^{\prime})n({\mathbf{r}}_{2}^{\prime})}\over{{r^{\prime}_{12}}}}\,d{\mathbf{r}}_{1}^{\prime}d{\mathbf{r}}_{2}^{\prime}}$ (2)
is the interionic interaction energy per ion, $r_{12}=|{\mathbf{r}}_{1}-{\mathbf{r}}_{2}|$, $n({\mathbf{r}}^{\prime})=\sum\nolimits_{k=1}^{N}{\delta({\mathbf{r}}^{\prime}-{{\mathbf{r}}_{k}})}$ is the ion number density, ${\mathbf{r}}_{k}$ is the radius-vector of the $k$th ion, and the integration is performed in the background volume $V_{R}$;
@@
$={1\over{2N}}\int\limits_{{V_{R}}}{{{r^{\prime}}^{2}}n({\mathbf{r}^{\prime}})}\,d{\mathbf{r}^{\prime}}-{5\over 2}{u_{c}}$
@@
is the background energy per ion. Equation (3) implies that the origin of the coordinate system coincides with the background center. Although $u$ converges at $N\to\infty$, $u_{p}$ and $u_{b}$ diverge as $N^{2/3}$; however, one can define renormalized quantities that will be termed the excess interatomic and ion–background electrostatic energy
@@
We will demonstrate convergence of $u_{p\mathrm{ex}}$ and $u_{b\mathrm{ex}}$ in the cases of weakly and strongly coupled systems. If the system is gaslike, and the ternary and higher-order ion correlations can be neglected, then
$\left\langle{n({{\mathbf{r}}_{1}})n({{\mathbf{r}}_{2}})}\right\rangle\simeq\left\langle{n(0)n(r_{12})}\right\rangle={\left({\displaystyle{3\over{4\pi}}}\right)^{2}}{f_{p}}(r_{12}),$ (6)
where brackets denote time averaging,
@@
is the radial distribution function, and $N_{r}$ is the average number of ions within a sphere of the radius $r$ around an ion. Similarly, $\left\langle{n({\mathbf{r}})}\right\rangle=\left\langle{n(r)}\right\rangle=(3/4\pi){f_{c}}(r)$, where
@@
$\left\langle{{u_{p\mathrm{ex}}}}\right\rangle$ $\simeq{1\over 2N}{\left({\displaystyle{3\over{4\pi}}}\right)^{2}}\iint\limits_{V_{R}}{\displaystyle{{f_{p}}(r_{1}){f_{p}}(r_{2})-1\over{r_{12}}}\,d{\mathbf{r}}_{1}d{\mathbf{r}}_{2}}$ (9)
@@
$\left\langle{{u_{b\mathrm{ex}}}}\right\rangle=\displaystyle{3\over{2N}}\int\limits_{0}^{R}{{r^{4}}\left[{{f_{c}}(r)-1}\right]\,dr}.$ (10)
@@
Thus, in the thermodynamic limit, we obtain an obvious result: all excess electrostatic energies of an ideal gas vanish:
@@
where $u_{\infty}=\mathop{\lim}\limits_{N\to\infty}\left\langle{u}\right\rangle$. Note that if we treat separately the system of ions in the field of background then the ionic potential energy $u_{i}={u_{p}}+{u_{b}}$ diverges as $N\to\infty$. However, the excess ionic potential energy $u_{i}+u_{c}={u_{p\mathrm{ex}}}+{u_{b\mathrm{ex}}}$, which coincides with the excess total electrostatic energy, does converge.
Consider the case $\Gamma\to\infty$, in which the ion sphere model (see, e.g., [2]) proved to be most effective for the Coulomb clusters [75, 73]. This model implies that the entire system is divided into spherical overlapping spheres of the radius $r_{s}$, in which single ions reside, and all interparticle and particle–background interaction can be reduced to the interaction of an ion with the compensating background of its sphere. Thus, the net force acting on an ion is simply ${\mathbf{f}}=-{\mathbf{r}}$, where the origin of the coordinate system is shifted to the center of an ion sphere and we omit subscripts due to identity of the ions. Based on the virial theorem [76] and an obvious equality for a harmonic oscillator $\left\langle{{r^{2}}}\right\rangle=\left\langle{{v^{2}}}\right\rangle=3/\Gamma$, we can write the ionic compressibility factor as [73]
${Z_{i}}=1+\displaystyle{\Gamma\over{3N}}\sum\limits_{i=1}^{N}{\left\langle{{\mathbf{rf}}}\right\rangle}=1-\displaystyle{\Gamma\over 3}\left\langle{{r^{2}}}\right\rangle=0,$ (13)
@@
We will use the virial theorem one more time to express the ion pressure and corresponding compressibility factor in terms of the ion subsystem potential and kinetic energy. However, if we assume that the charge of neutralizing background of BOCP is fixed, $r_{s}=\mathrm{const}$, then $u_{p}$ (Eq.2) is a homogeneous function of order 1-1; $u_{b}+5u_{c}/2$, of order 2 (Eq. (3)); and $u_{c}$, of order 0. Bearing in mind that if the function $\varphi$ is a sum of homogeneous functions of order $k$, $\varphi(x)=\sum\nolimits_{k}{{\psi_{k}}(x)}$, then
${{d\varphi}\over{dx}}=\sum\limits_{k}{\displaystyle{{k{\psi_{k}}(x)}\over x}}.$ (14)
Thus, one can modify the virial theorem for the ions in the field of background as
@@
Thus, we have demonstrated that the quantities (5) have thermodynamic limits at low and high $\Gamma$, which is indicative of the fact that these limits exist at arbitrary $\Gamma$ defining the ionic pressure for the bulk OCP.
Since an objective of this paper is investigation of the OCP properties, the size dependence $u(\Gamma,N)$ (1) is of interest. As it was demonstrated in [73], all quantities derived from the BOCP energy characteristics depend on NN very weakly, therefore, we adopt $\ln N$ as an argument. The simplest and most general form of the size dependence is then the expansion in powers of shifted $\ln N$, where we retain only two terms:
@@
Here, $u_{\infty}(\Gamma)$, $c_{1}$, and $c_{2}$ are the parameters to be defined from fitting the MD simulation results (Sec. III.2).